n small drops of same size are charged to V volts each. If they coalesce to form a single large drop, then its work done will be?

Grade:12

1 Answers

Vikas TU
6 years ago
Give V a chance to be the potential contrast. Give V of 1 a chance to drop = kq/r. 
In the event that the volume of n drops = n 4/3πr3 
Volume of greater drop =4/3πR3 
At that point from the question 
Volume of n drops = Volume of greater drop 
n 4/3πr3 =4/3πR3 
(n1/3r) = R 
As charge stays saved 
As 1 drop has charge q so the n drops will have nq charge. 
Work done(Potential distinction (V)) = knq/(n1/3r) 
= n2/3 kq/r 
= n2/3 V